OpenStudy (anonymous):

A toy rocket is launched vertically from ground level (y = 0 m), at time t=0 s. The rocket engine provides constant upward acceleration during the burning phase. At the instant of engine burnout, the rocket has risen to 68 m and acquired a velocity of 30 m/s. The rocket continues to rise in unpowered flight, reaches maximum height, and falls back to the ground.

OpenStudy (anonymous):

1.The time interval, during which the rocket engine provides upward acceleration, is closest to 5.0 s 3.0 s 4.0 s 4.5 s 3.5 s 2.The upward acceleration, in m/s2, of the rocket during the burn phase is closest to 6.0 5.8 6.4 6.2 6.6 3.The maximum height, in meters, reached by the rocket is closest to 125 105 115 120 110 4.The time interval, during which the rocket is in unpowered flight, is closest to 7.0 s 7.5 s 8.0 s 6.5 s 6.0 s 5.The speed of the rocket, in m/s, upon impact on the ground is closest to 37 52 47 57 42

OpenStudy (theeric):

So, you have \(y_0\), \(v_0\), \(t_0\), and the \(y\) and \(v\) at which the rocket stops accelerating. Do you have an equation in mind?

OpenStudy (anonymous):

wait its not vf^2=vi^2+2(a)(d) is it?

OpenStudy (theeric):

It definitely is!!

OpenStudy (theeric):

Wait, not that one.. There's no time, t! One sec.

OpenStudy (theeric):

\[x−x_0 =\frac{v_0 +v_2}{ 2}\ t\] from @Festinger

OpenStudy (anonymous):

whats x= again?

OpenStudy (theeric):

Make that a \(y\), since we're dealing with the vertical :)

OpenStudy (theeric):

\(y\) is the height at the end of the interval.

OpenStudy (anonymous):

so 68-0= (0+30/2)(t) which ---> 4.53=t

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

ok can you help me with the next part of the question

OpenStudy (anonymous):

... and how do i differentiate between m/s and m/s2

OpenStudy (theeric):

That's what I got! So now you have \(\Delta t\), the time elapsed while the rocket was accelerating upwards. There's a handy equation that is pretty much the definition of acceleration put into math. Do you know it?

OpenStudy (theeric):

It was actually equation (2) from Festinger. \[a=\frac{\Delta v}{\Delta t}\]

OpenStudy (anonymous):

haha i actually knew that one!

OpenStudy (anonymous):

so 30-0/4.5-0 which ---> a=6.6 is that right?

OpenStudy (anonymous):

ok next part

OpenStudy (theeric):

Again, got the same thing! Sorry, I'm multitasking. For this one did you have an equation in mind?

OpenStudy (anonymous):

x=x0+v0t+1/2at^2 x−x0=v0+v2t

OpenStudy (anonymous):

one of those right?

OpenStudy (theeric):

And m/s is velocity, while \(m/s^2\) is acceleration. You have to make sure you're thinking of the right variables for this new interval, which is after the rocket boosted up. Now you don't necessarily know the time interval until it reaches the top. But you do know it stops at the top before heading down, so the final velocity is 0. To start, you know you're up 68m, and you have a velocity of \(30\ m/s\) to start with and \(0\ m/s\) at the end. You use \(-9.8\ m/s^2\) for gravitational acceleration, right?

OpenStudy (theeric):

Or \(-10\ m/s^2\)?

OpenStudy (anonymous):

right

OpenStudy (theeric):

For this second interval, you can use \(v^2 = v_0^2+2ad\) to find the distance while the rocket was not boosting. Then you can add it to the 68m you were up already.

OpenStudy (anonymous):

i got 45.91+68 to get 114 so about 115 correct?

OpenStudy (theeric):

Yeah! I got 45 when using \(10\ m/s^2\), and same as you with \(9.8\ m/s^2\), so yeah!

OpenStudy (anonymous):

perfect! how bout the next question

OpenStudy (theeric):

Well, the first part of unpowered flight is what we just worked with. The initial \(v\) was \(30\ m/s\), the final \(v\) was \(0\). The acceleration was \(-10\ m/s\), so you can use the definition of acceleration again!

OpenStudy (anonymous):

but that didnt work... i got 3, wouldnt i need at east one time (initial or final)

OpenStudy (anonymous):

what did you get?

OpenStudy (theeric):

Yeah, that interval is about 3s. But then there's the entire fall to go yet! You know how high you are (\(d\)), and the acceleration from gravity (\(-10\ m/s^2\)), and the velocity you start with at the top. You need the time!

OpenStudy (theeric):

I sort of want to find #5 first... To get the final velocity. But there's probably an equation to go on with #4!

OpenStudy (anonymous):

lets do number 5 first then

OpenStudy (theeric):

I agree.

OpenStudy (anonymous):

i honestly have no clue how you do number 4 or 5

OpenStudy (theeric):

Haha, #5, then.

OpenStudy (theeric):

So, let me draw a picture. They usually help. It'll be sloppy, but quick. I'll draw only the fall.

OpenStudy (theeric):



OpenStudy (theeric):

\[v^2 = v_0^2+2ad\]

OpenStudy (anonymous):

whats d=

OpenStudy (anonymous):

115 or 68?

OpenStudy (anonymous):

115 right?

OpenStudy (theeric):

115, yeah!

OpenStudy (theeric):

That was our max height...

OpenStudy (anonymous):

would we want to use 9.8 sense it is asking in m/s

OpenStudy (theeric):

\[v = \sqrt{v_0^2+2ad}\]

OpenStudy (theeric):

\(a\) is about \(-10\ m/s^2\), and \(v\) will come out to be in \(m/s\).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 47.95 so about 47 right?

OpenStudy (theeric):

I got the same! Back to #4!

OpenStudy (theeric):

We found the freefall going up to be 3s. Now we have...And the handy definition of acceleration says that the time for the descent is...

OpenStudy (theeric):

\[a=\frac {v-v_0}{\Delta t}\]

OpenStudy (anonymous):

i got 4.7 thats not rght

OpenStudy (theeric):

But that was only the freefall descent. There was also the freefall ascent. You see where I'm going?

OpenStudy (anonymous):

you add them? idk

OpenStudy (theeric):

Take a guess.

OpenStudy (anonymous):

is it 7.5?

OpenStudy (theeric):

That's what I got at first, but it was sort of between answers. So I used the real values. The time up was 30/9.8 and the time down was about 48/9.8, and I added them to get about 8s.

OpenStudy (theeric):

So, I think that's the answer it should want...

OpenStudy (anonymous):

yea great thank you so much! i really appreciate it, and im glad you actually taught me how to do it instead of just giving me the answers... haave a good night and maybe you can help me again in the future. Thanks!

OpenStudy (theeric):

It's about 7.75 with \(a=10\ m/s^2\) though....

OpenStudy (anonymous):

8 worked perfectly

OpenStudy (theeric):

I'd like to help if possible! And I wish those answers weren't so close together! Did you get them all right?

OpenStudy (anonymous):

yup! 100% (it looks like there werent any point deductions for late answers either :) )

OpenStudy (anonymous):

im going now but thank you again!

OpenStudy (theeric):

You're very welcome! And you did well! :)

OpenStudy (theeric):

Glad it all worked out.